Circular No. 6445

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995N IN MCG –2-38-017 W. H. G. Lewin, Massachusetts Institute of Technology; and H.-U.

Zimmermann and B. Aschenbach, Max-Planck-Institut für Extraterrestrische Physik, report: “On July 23, the ROSAT satellite’s HRI detector was pointed toward SN 1995N for a total exposure of 1331 s. About 3′′ off the optical position (IAUC 6170, 6386), which is well within the ROSAT error box, a pointlike x-ray source was detected with very high significance. During the ROSAT All-Sky Survey in 1990/1991, no x-ray source of similar strength was observed at that position; we therefore identify this x-ray source with SN 1995N. The observed rate was 0.0075 ± 0.0025 count/s. Assuming a galactic hydrogen column density of N_{H} = 0.8 × 10^{21} cm^{−2}, a thermal spectrum with kT ∼ 1 keV, and the distance to the host galaxy as 24 Mpc, the observed countrate corresponds to an (unabsorbed) x-ray luminosity of 2–3 × 10^{40} erg/s in the ROSAT band (0.1–2.4 keV).”

RX J0720.4–3125 F. Haberl and W. Pietsch, Max-Planck-Institut für Extraterrestrische

Physik, Garching; C. Motch, Observatoire de Strasbourg; and D. A. H. Buckley, South African Astronomical Observatory (SAAO), Cape Town, communicate: “We have discovered what appears to be an isolated, xray-pulsating neutron star that is accreting matter from the interstellar medium. The object was found in the ROSAT All-Sky Survey as a soft x-ray source with 1.6 counts/s. A ROSAT observation using the Position Sensitive Proportional Counter in 1993 revealed a blackbody-like x-ray spectrum with kT = 75–83 eV, attenuated by little absorption with a hydrogen column density of 1.3 × 10^{20} cm^{−2}. The best determined position of α = 7^{h}20^{m}25^{s}.04, δ = −31^{o}25′48′′.4 (equinox 2000.0; uncertainty ± 2′′ using a nearby optically identified source as reference) was derived from an observation with the ROSAT High Resolution Imager on 1996 May 7. In all pointed ROSAT observations, a periodic modulation of the x-ray flux is detected with a period of 8.38 s. No long-term variations of more than ±10 percent on time scales of years between the ROSAT observations are seen in the source intensity, and archival Einstein IPC and EXOSAT LE detections are consistent with this. CCD images obtained at the SAAO failed to detect an optical counterpart to a limiting magnitude of V ∼ 21.2. We encourage multi-wavelength observations to confirm this candidate.”

1996 August 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green